# FinTruthQA: A Benchmark Dataset for Evaluating the Quality of Financial Information Disclosure

### v1

@@ 6 Conclusion
    For all authors…
    1. (a)
      Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes]
    2. (b)
      Did you describe the limitations of your work? [Yes] See Section 5. Limitation.
    3. (c)
      Did you discuss any potential negative societal impacts of your work? [Yes] See Supplement Section A.
    4. (d)
      Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes] More discussion can be found in Supplement Section A.
@@
    If you are including theoretical results…
    1. (a)
      Did you state the full set of assumptions of all theoretical results? [N/A] This study aims for the construction and evaluation of a new benchmark. No theoretical assumptions.
    2. (b)
      Did you include complete proofs of all theoretical results? [N/A] This study aims for the construction and evaluation of a new benchmark. No theoretical assumptions.
@@
    If you ran experiments (e.g. for benchmarks)…
    1. (a)
      Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [Yes] All the datasets, benchmarks and code are available at https://github.com/bethxx99/FinTruthQA
    2. (b)
      Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [Yes] See Section 3.5 for data splits and Supplement Section E for hyperparameter choices.
    3. (c)
      Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? [Yes] The detailed results can be found in the supplementary material.
    4. (d)
      Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes] See Supplement Section E.
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
    1. (a)
      If your work uses existing assets, did you cite the creators? [Yes] We used code from several models in our benchmarks, all the sources were properly cited in this paper.
    2. (b)
      Did you mention the license of the assets? [Yes] Our code and data are open access and license information is listed in GitHub.
    3. (c)
      Did you include any new assets either in the supplemental material or as a URL? [Yes] See Supplementary materials.
    4. (d)
      Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [N/A] This data used in this study are open-available.
    5. (e)
      Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [Yes] See Supplement Section A.
  5. 5.
    If you used crowdsourcing or conducted research with human subjects…
    1. (a)
      Did you include the full text of instructions given to participants and screenshots, if applicable? [N/A] This study does not involve crowdsourcing and human subjects.
    2. (b)
      Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [N/A] This study does not involve crowdsourcing and human subjects.
    3. (c)
      Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [N/A] This study does not involve crowdsourcing and human subjects.

### v2

@@ Appendix B Annotation process
    Pre-annotation: Prior to the formal annotation, multiple rounds of pre-annotation were conducted to refine the annotation standards. Each round of pre-annotation involved 100 samples. Six annotators were trained and provided with the evaluation criteria and examples. After becoming familiar with the criteria, they rated the provided samples. After each round, the inter-annotator agreement was calculated. Based on the inconsistent parts, the relevant annotation rules were revised according to the discussion results. This process continued for a total of 5 rounds. Eventually, four annotators with the highest consistency ratings were selected to move onto the formal annotation stage.
@@
    Formal annotation: The formal annotation process also used a multi-round iterative mode, with each Q&A sample annotated by two annotators, A and B. Any uncertainties or disagreements between annotators were recorded in Tencent documents. A and B initially address their disagreement, and the project leader quantified the consistency of the annotation results, identifying and documenting the areas where inconsistencies and uncertainties were found. These issues were then discussed with experts to reach a consensus. After the discussion, A was responsible for modifying the annotation to form the final version of the annotation file.
@@
    Data verification and integration: After completing the formal annotation, the project leader checked whether the annotated data conforms to the latest version of the annotation criteria, as well as whether there were missing scores and other issues. Once confirmed, all annotated data was integrated, and key information about the data set, such as sample size and score distribution in each dimension, is summarized.
Appendix C Examples of mislabeled data points and quality control
{CJK}
UTF8gbsn
Table 6: Sample data points that were mislabeled.
Question Answer Label
您好,请问大股东股权质押的平仓线是多少?$谢$ 尊敬的投资者,公司未收到大股东出现平仓情况的通知。感谢您对贵人鸟的关注。 1
Hello, may I ask what is the liquidation line for the major shareholder’s share pledge? Thank you. Dear investors, the company has not received any notification regarding the occurrence of forced liquidation of major shareholders. Thank you for your attention to Guirenniao.
公司领导为什么不学李东生,在大会上说公司估值太低?常年二三元的股价,有损公司形象? 您好,感谢您对公司的关注。 3
Why don’t company leaders learn from Dongsheng Li and say at the conference that the company is undervalued? The stock price has been around two to three yuan for years, damaging the company’s image? Hello, thank you for your interest in the company.
董秘你好,云海金属经过多年发展已经成长为一个集科研,垂直产业一体化的公司,不仅仅是一个单一有色金属的公司,建议公司更名,一是拓展自身的发展愿景,二是让更多的投资者知道了解云海,融资助力云海更上一层楼。建议更名“云海科技”,或者“云海创新”,改名绝不是蹭热度,混时尚,而是极具战略意义的宣言。希望公司采纳,谢谢。 好的,谢谢您的建议。 1
Dear Secretary, Yunhai Metals has grown into a company that integrates scientific research and vertical industry integration after years of development. It is no longer just a company that deals with non-ferrous metals. It is suggested that the company change its name to expand its own development vision and let more investors know about Yunhai, aiding in financing and helping Yunhai reach a new level. The suggested names are ”Yunhai Technology” or ”Yunhai Innovation”. Changing the name is not about chasing trends or blending into the fashion scene but is a declaration of significant strategic meaning. I hope the company will consider this suggestion. Thank you. Okay, thank you for your suggestion.
Sample data points that were mislabelled are presented in Table 6. Our approach of quality control involved training the FinBERT model on the complete dataset for a limited duration of 5 epochs, a precautionary measure to prevent overfitting. Subsequently, this trained model was utilized to predict the entire dataset. The data points where the model’s predictions were inconsistent with the labels annotated by experts were extracted and sorted in ascending order according to their predicted probabilities. These sorted data points were then presented to experts for further evaluation. This method aligns with the CL framework, which emphasizes the identification and correction of label errors in a structured manner (Northcutt et al., 2021b). By utilizing predicted probabilities as indicators of prediction confidence, our method facilitated a focused review of the most challenging samples for the model. This strategy allowed experts to concentrate on instances where the model faced difficulty in classification. The experts could then revisit the labeling instructions and make necessary corrections. The revised annotations were then utilized to retrain the model, contributing to an refined dataset and subsequently improved model performance.
The process was conducted for the answer relevance task over two rounds. In the first round, 201 QA pairs were identified where the model predicted incorrectly. We had the annotators re-evaluate and correct the labels for 91 of them. In the second round, we identified an additional 151 data that required re-evaluation, resulting in 38 corrected labels. The distribution of labels for this task presented in this paper was obtained after the quality control process.
Appendix D Pre-trained language model
Table 7 provides links to specific PLMs we used in this study.
Table 7: Links for different PLMs used in our experiments.
Appendix E Hyperparameter settings
Table 8: Hyperparameters for BERT-based models. The embedding size is 768 for the base model and 1024 for the large model. Other hyperparameters are shared across both configurations.
Parameter Value Parameter Value
learning rate 2e-5 embedding size (base) 768
dropout rate 0.1 embedding size (Large) 1024
number of epochs 30 batch size 32
lr decay 0.01 max length 256
Table 8 presents the hyperparameter settings of our BERT-based model, which were kept fixed across all experiments except from the embedding size. These hyperparameters were initialized based on prior work in the literature. During our initial experiments, we also fine-tuned these parameters to strike a suitable balance between model complexity and performance using the validation set from the preprocessed dataset. In particular, the embedding size for the BERT-base model was set to 768, while it was set to 1024 for BERT-large model. AdamW (Loshchilov & Hutter, 2018) was used as the optimizer in all experiments as we found that AdamW exhibited better stability and faster convergence compared to stochastic gradient descent. The experiments were conducted using two NVIDIA RTX 3090 GPUs.
Appendix F Example prompt for LLM-based experiments
The prompts of LLMs are shown in table LABEL:prp, which include two versions: ”Instructions to Score Answers” and ”Extended Prompts Requiring Analysis.” The former instructs the model to directly output relevance scores, while the latter requires additional analysis output. The prompts provided to LLMs were designed as a classification task. We instructed the model to assign a score of 1, 2, or 3, with each score corresponding to a specific level of relevance. This design aligns with a typical classification task, where the model is required to categorize the answers based on predefined criteria. The scoring system was explicitly explained to the model, and the model’s outputs were treated as categorical classifications, thus ensuring that the evaluation remained consistent with the other tasks.
{CJK}
UTF8gbsn
Table 9: Prompt for LLM-based experiments. The text in purple instructs the model to provide an analysis of its decision-making process alongside the relevance scores. This instruction was only included in the prompts for the second experiment to gather insights into LLMs’ rationale for scoring answer relevance.
Instructions to Score Answers
## Instructions
假设您是一名经验丰富的金融信息分析师,您需要从信息披露的角度来判断Q&A数据的质量。请您根据以下评分标准,对每个问题的回答进行判断并给予相应的分数,并按照’回答相关性评分为:…’的格式输出评分:
(Translation: Assume you are an experienced financial information analyst. You need to evaluate the quality of Q&A data from the perspective of information disclosure. Please judge each answer based on the following scoring criteria, assign a corresponding score, and output the score in the format: ”The relevance score of the answer is: …”.)
## Evaluation Criteria
评分标准:
1分:回答与提问完全不相关(完全不相关指所问问题一个也没有回答)
2分:部分相关(部分相关指回答了部分问题)或难以判断
3分:回答与提问完全相关(完全相关指回答了提问中的全部问题)
说明:回答相关性要求公司对投资者的提问一一进行正面回复。所谓相关,指回答正面回答了提问者的问题,不回答或者是回答与问题主题有关但并未正面回答提问都为无关。正面回复了提问者的所有问题,则为完全相关,得3分。所有的问题都没有正面回复,则为完全不相关,得1分。正面回答了提问中的部分问题,则为部分相关,部分相关或是难以判断的回答得2分。如果一个回答能回答多个问题,也为全部相关,得3分;如果提问中提到的问题与公司无关,公司只需要否定,不需要回答后续问题。
(Translation: Scoring criteria:
1 point: The answer is completely unrelated to the question (completely unrelated means none of the questions were answered).
2 points: Partially related (partially related means some questions were answered) or difficult to judge.
3 points: The answer is fully related to the question (fully related means all questions were answered).)
## Input
问题(Question): {question}
回答(Answer): {answer}
Extended Prompts Requiring Analysis
## Instructions
假设您是一名经验丰富的金融信息分析师,您需要从信息披露的角度来判断Q&A数据的质量。请您根据以下评分标准,对每个问题的回答进行判断并给予相应的分数,并按照’分析过程为:…因此,回答相关性评分为:…’的格式输出评分:
(Translation: Assume you are an experienced financial information analyst. You need to evaluate the quality of Q&A data from the perspective of information disclosure. Please judge each answer based on the following scoring criteria, assign a corresponding score, and output the score in the format: ”Analysis process: …, therefore, the relevance score of the answer is: …”.)
## Evaluation Criteria
评分标准:
1分:回答与提问完全不相关(完全不相关指所问问题一个也没有回答)
2分:部分相关(部分相关指回答了部分问题)或难以判断
3分:回答与提问完全相关(完全相关指回答了提问中的全部问题)
说明:回答相关性要求公司对投资者的提问一一进行正面回复。所谓相关,指回答正面回答了提问者的问题,不回答或者是回答与问题主题有关但并未正面回答提问都为无关。正面回复了提问者的所有问题,则为完全相关,得3分。所有的问题都没有正面回复,则为完全不相关,得1分。正面回答了提问中的部分问题,则为部分相关,部分相关或是难以判断的回答得2分。如果一个回答能回答多个问题,也为全部相关,得3分;如果提问中提到的问题与公司无关,公司只需要否定,不需要回答后续问题。
(Translation: Scoring criteria:
1 point: The answer is completely unrelated to the question (completely unrelated means none of the questions were answered).
2 points: Partially related (partially related means some questions were answered) or difficult to judge.
3 points: The answer is fully related to the question (fully related means all questions were answered).)
## Input
问题(Question): {question}
回答(Answer): {answer}
Appendix G Data diversity assessment
The questions of our dataset cover a wide range of topics, including project and investment inquiries (e.g., questions about specific project, investment impacts and progress), financial and stock performance (e.g., inquiries about stock price movements, shareholding structures and potential market manipulation), operational and production issues (e.g., questions related to production plans and operational status during the pandemic), market and competitive analysis (e.g., concerns about market share and competition in specific sectors), raw material and supply chain (e.g. inquiries about material usage and supply chain issues), etc. The questions also vary in depth, ranging from surface-level queries about basic facts (e.g., number of shareholders) to more detailed questions requiring in-depth responses, such as the implications of specific projects on future earnings or strategic decisions regarding market expansion. These examples indicate that our dataset is not limited to a single type of question but instead captures a broad spectrum of investor concerns, ranging from financial and stock performance to operational details and market competition.
{CJK}
UTF8gbsn
Case 1: About project and investment inquiries
问题(Question):董秘您好,请问大连项目是否已经投产,会对今年业绩产生积极影响吗?(Dear Secretary of the Board, could you please confirm if the Dalian project has already commenced production? Will it have a positive impact on this year’s performance?)
答案(Answer):尊敬的投资者,目前大连项目处于试运行阶段,试运行阶段会产生部分的收益,预计本年度不会对公司业绩产生较大影响。谢谢!(Dear Investors, the Dalian project is currently in the trial operation phase. This phase will generate some revenue, but it is not expected to have a significant impact on the company’s performance this year. Thank you!)
Case 2: About financial and stock performance
问题(Question):火炬是否经营正常,股价最近很低迷,是否存在恶意卖空,陶瓷材料新业务预计什么时候可以产生投资收益,谢谢!(Is the Torch project operating normally? The stock price has been sluggish recently—could there be malicious short selling? When is the new ceramics materials business expected to generate investment returns? Thank you!)
答案(Answer):您好!公司经营状况一切正常。另外定增项目已完成三条(总共六条)材料生产线的安装调试工作,已进入试生产阶段,待相关鉴定及验证完成后,可释放5吨/年的产能,形成新的利润增长点。谢谢关注。(Hello! The company’s operations are running smoothly. Additionally, three out of six production lines for the materials in the private placement project have completed installation and commissioning and have entered the trial production stage. Once the relevant assessments and verifications are completed, a production capacity of 5 tons per year will be released, creating a new profit growth point. Thank you for your attention.)
Case 3: About market and competitive analysis
问题(Question):你好,近期菜油和玉米期货大涨,能否介绍下公司油脂业务方面的情况和市场占有率,以及玉米大宗货物交易方面的业务情况和中国玉米网的市场地位。(Hello recently, there has been a significant increase in the futures prices of rapeseed oil and corn. Could you please provide an overview of the company’s operations in the oils and fats business, including its market share? Additionally, could you explain the company’s business activities in bulk corn trading and the market position of China Corn Network?)
答案(Answer):尊敬的投资者您好。公司油脂业务经营主体为厦门东方银祥油脂有限公司和上海东瑞贸易有限公司,主营产品为豆油(一级豆油,四级豆油、大豆原油)、菜油(一级菜油,四级菜油,拍卖油)、豆粕、菜籽粕、磷脂油等,产品销售覆盖四川、重庆、贵州、广西、安徽等省市。公司玉米收购销售业务经营模式主要包括自营业务、港口中转业务和政策性粮贸易业务。公司旗下中国玉米网(中国玉米供应链服务平台)是国内玉米领域最大的信息集成平台,平台积极促进贸易对接,使玉米产业链上下游企业在线上更便捷地完成交易,并提供全面高效的线上数据信息服务。该平台自2020年2月上线至2020年6月末累计成交总额为15.12亿元。谢谢。(Dear Investor, thank you for your inquiry. The company’s oil business is primarily managed by Xiamen Dongfang Yixiang Oil Co., Ltd. and Shanghai Dongrui Trading Co., Ltd. The main products include soybean oil (first-grade soybean oil, fourth-grade soybean oil, crude soybean oil), rapeseed oil (first-grade rapeseed oil, fourth-grade rapeseed oil, auction oil), soybean meal, rapeseed meal, lecithin oil, and others. These products are sold in various provinces and cities, including Sichuan, Chongqing, Guizhou, Guangxi, and Anhui. The company’s corn procurement and sales business operates through various models, including self-operated business, port transshipment business, and policy-based grain trade business. The China Corn Network (China Corn Supply Chain Service Platform), a subsidiary of the company, is the largest information integration platform in the domestic corn industry. The platform actively promotes trade matching, enabling upstream and downstream enterprises in the corn industry chain to conduct transactions more conveniently online while providing comprehensive and efficient online data and information services. From its launch in February 2020 to the end of June 2020, the platform achieved a total transaction value of 1.512 billion yuan. Thank you.)
The variety ensures that the dataset is robust and capable of supporting comprehensive analysis across multiple dimensions of corporate communication and financial disclosure quality. This diversity enhances the value of our dataset, making it a valuable resource for understanding the different facets of company-investor interactions.